f(x)=(3x^(3)+5)*e^(2x^(2))


1
Expert's answer
2021-09-28T09:19:45-0400

"\text{Given the function $f(x)=(3x^{3}+5)*e^{2x^{2}}$}\\
\text{We find the derivative of f(x) using product rule}\\
\text{Let $u = 3x^{3}+5$ and $v = e^{2x^{2}}$ }\\
\text{Hence, $f'(x) = e^{2x^2} \cdot 9x^2 + (3x^3+5) \cdot 4xe^{2x^2}$}\\
f'(x) = e^{2x^2}(9x^2+12x^4+20x)"